OpenStudy (anonymous):

The offical height-to-width ratio of the United States flag is 1 to 9. If a United States flag is 9.5 feet wide how high should it be?

OpenStudy (anonymous):

height-to-width ratio = height/width height/width = 1/9 height/9.5 = 1/9 height = 9.5/9 height = 1.056